The newly constructed bridge in your town is a steel arch bridge 600m in length. How much does the total length of the roadway decking change between temperature extremes of 30⁰C and 50⁰C?


1
Expert's answer
2021-05-04T12:07:36-0400

"\frac{ΔL}{L_0} = αΔT \\
ΔL = αL_0ΔT \\
ΔL = 11 \times 10^{-6} \times 600 (50 -30) \\
= 0.132 \;m \\
= 13.2 \;cm"